Suppose the demand curve for a product is given by Q = 10 − 2P + PS, where P is the price of the product and PS is the price of a substitute good. The price of the substitute good is $2.00
1
Expert's answer
2020-12-10T14:29:45-0500

Suppose the demand curve for a product is given by Q = 10 − 2P + PS, where P is the price of the product and PS is the price of a substitute good. The price of the substitute good is $2.00. Suppose P =$1.00. What is the price elasticity of demand?

"Q = 10 - 2P + P_s \\

\frac{dQ}{dP} = -2 \\

P = 1 \\

P_s = 2 \\

Q = 10 - 2 \times 1 + 2 \\

Q = 10 \; units"

Price elasticity of demand "= \frac{dQ}{dP} \times \frac{P}{Q} = -2 \times \frac{1}{10} = -0.2"

Price elasticity of demand = -0.2